OpenStudy (anonymous):

What is the output of f(x)=-2-3x when x=-4/3?

OpenStudy (anonymous):

plug in for x, what do you get?

OpenStudy (anonymous):

f(-4/3)=-2-3(-4/3)

OpenStudy (anonymous):

correct, now calculate it

OpenStudy (anonymous):

Is it -6?

OpenStudy (anonymous):

nope, you made a small error

OpenStudy (anonymous):

Positive 6? I think I messed it up with the negatives.

OpenStudy (anonymous):

ok -2 -3*(-4/3) what is 3*(-4/3)?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

ok -2 - (-4) = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

That's the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

was it hard?

OpenStudy (anonymous):

No, not as hard as I thought it would be! Thank you very much.

OpenStudy (anonymous):

no problem, just pay attention to the negatives and you will be fine :)